OpenStudy (anonymous):

6sqrt(x)(x^3-8sqrt(x)+7) findf'(9)=

OpenStudy (anonymous):

\[f(x)=6\sqrt{x}(x^3-8\sqrt{x}+7)\]

OpenStudy (anonymous):

@stamp

OpenStudy (stamp):

Take the derivative. Distribute the \[6\sqrt x \]to obtain\[6x^{7/2}-48x+42\sqrt x\]Take the derivative of each term. Solution: http://www.wolframalpha.com/input/?i=derivative+of+6x%5E.5%28x%5E3-8x%5E.5%2B7%29 Then evaluate f'(x) for f'(9).